OpenStudy (anonymous):

Suppose an airplane climbs 15 feet for every 40 feet it moves forward. What is the slope of this airplane's ascent?

OpenStudy (anonymous):

problems like these i like to draw out, but this one is easy

OpenStudy (anonymous):

this is a linear function where altitude is a function of distance

OpenStudy (anonymous):

so altitude is your y axis and distance is you x axis

OpenStudy (anonymous):

visually speaking because draw is broken

OpenStudy (anonymous):

essentially we have an ordered pair (x,y)

OpenStudy (anonymous):

(40,15)

OpenStudy (anonymous):

what should x be when we have (x,30)

OpenStudy (anonymous):

i dont know what

OpenStudy (anonymous):

well it climbed 15 feet

OpenStudy (anonymous):

and every 15 feet you go up 40

OpenStudy (anonymous):

so (80,30) is your next ordered pair

OpenStudy (anonymous):

do you know how to find slope?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

when you have two ordered pairs

OpenStudy (anonymous):

howd u get 80

OpenStudy (anonymous):

we started at 40, so i just doubled it

OpenStudy (anonymous):

but its gonna turn out a fraction

OpenStudy (anonymous):

slope isn't always a perfectly round number

OpenStudy (anonymous):

you can have a slope of 1 or a slope of 1/39

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

so wat do i do now

OpenStudy (anonymous):

you have two ordered pairs (40,15), (80,30)

OpenStudy (anonymous):

so what do we do now

OpenStudy (anonymous):

the plane is ascending the skies, so this is a positive slope by the way

OpenStudy (anonymous):

do you know the slope formula

OpenStudy (anonymous):

y=mx+b

OpenStudy (anonymous):

that is point slope form!

OpenStudy (anonymous):

we are looking for m in that equation

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

m=(y2-y1)/(x2-x1)

OpenStudy (anonymous):

so m= 30-15/80-40

OpenStudy (anonymous):

15/40

OpenStudy (anonymous):

which can be simplified

OpenStudy (anonymous):

3/8

OpenStudy (anonymous):

oh yea i kno what it is

OpenStudy (anonymous):

gratis good luck on hw peace

OpenStudy (anonymous):

bed time

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

whats the answer